Find the area using definite integration


1.) y = 3x2 from x = 1 to x = 2


2.) xy = -1 from x = 1 to x = 2


3.) y = 4 - x2 from x = -3 to x = 3


1
Expert's answer
2021-05-07T10:08:59-0400

1) y = 3x2 from x = 1 to x = 2

"\int_{1}^23x^2dx=3\int_{1}^2x^2dx=3[\frac{x^{2+1}}{2+1}]|_1^2=3[\frac{x^{3}}{3}]|_1^2=x^3|_1^2=2^3-1^3 =8 - 1=7"

2) xy = -1 from x = 1 to x = 2

"xy=-1\implies y = -\frac{1}{x}"

"\int_1^2-\frac{dx}{x}=-[ln|x|]|_1^2=-(ln(2)-ln(1)) =-ln(2)"

3) y = 4 - x2 from x = -3 to x = 3

"\int_{-3}^3(4-x^2)dx=\int_{-3}^34dx-\int_{-3}^3x^2dx=4[x]|_{-3}^3-[\frac{x^3}{3}]|_{-3}^3=4(3-(-3))-(\frac{27-(-27)}{3})=12+12-(9+9)=24-18=6"